two kilograms of gas undergo a reversible non-flow process at 350 kPa while the volume increases from 0.4734 m^3. find the work done, kj/kg
1
Expert's answer
2020-03-18T10:25:14-0400

Heating of Fluid in a cylinder at Constant pressure {Isobaric Process}. Then work which done


"A=\frac{p\cdot \Delta V}{m}"

Calculating


"A=\frac{350\cdot 0.4734}{2}=82.845 (kJ/kg)"